# DiffMesh: A Motion-aware Diffusion-like Framework for Human Mesh Recovery from Videos

### v3

@@ 3.3 DiffMesh
Video-based Method Venue Number of frames MPVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ MPVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$
VIBE [20] CVPR 2020 16 - 65.6 41.4 99.1 82.9 51.9
TCMR [6] CVPR 2021 16 - 62.3 41.1 102.9 86.5 52.7
MAED [49] ICCV 2021 16 84.1 60.4 38.3 93.3 79.0 45.7
MPS-Net [50] CVPR 2022 16 - 69.4 47.4 99.7 84.3 52.1
SmoothNet [56] ECCV 2022 32 - 67.5 46.3 - 86.7 52.7
GLoT [42] CVPR 2023 16 - 67.0 46.3 96.3 80.7 50.6
MotionBERT [64] ICCV 2023 16 65.5 53.8 34.9 88.1 76.9 47.2
DiffMesh (Ours) 16 64.2 52.5 33.5 86.4 75.7 45.6
@@
DND [25] with HybrIK [24] ECCV 2022 32 - 52.5 35.5 88.6 73.7 42.7
MotionBERT [64] with HybrIK [24] ICCV 2023 16 52.6 43.1 27.8 79.4 68.8 40.6
DiffMesh (Ours) with HybrIK [24] 16 52.1 42.4 27.7 78.4 67.8 40.1
Table 2: Performance comparison with SOTA video-based methods on Human3.6M and 3DPW datasets. 3DPW Training set is used during training. The best results are marked as bold and the second-best results are marked as underlined.
3.4 Loss Functions
During training, we supervise the SMPL parameters ($\Theta_{i}$ and $\beta_{i}$) where $i\in\left\{1,\dots f\right\}$ with the ground-truth SMPL parameters of each frame using the $L2$ loss. Following [20, 6], the 3D joint coordinates are regressed by forwarding the SMPL parameters to the SMPL human body model. Furthermore, as mentioned in Sec. 3.3, Mean Squared Error (MSE) loss is applied between the conditional feature $c_{i}$ and the predicted conditional features $\hat{c_{i}}$. More detailed information is included in the supplementary Sec. B.
4 Experiments
@@ 4.1 Implementation Details
For fair comparisons, we adhere to the standard implementation details commonly employed in the video-based HMR task, consistent with prior works such as [20, 6, 50, 64]. We set the length of the input sequence $f$ to 16. The number of steps $N=30$. We implemented our approach using PyTorch [36] on a single NVIDIA A5000 GPU. The optimization of weights is conducted using the Adam optimizer [19], with an initial learning rate of $5e^{-6}$ and weight decay set to 0.98. During the training phase, we utilize a batch size of 64, and the training process spans 60 epochs. More detailed information about the conditional feature generation part is included in the supplementary Sec. B.
@@ 4.3 Comparison with State-of-the-art Methods
We compare our DiffMesh with previous SOTA video-based methods on Human3.6M and 3DPW datasets (results on MPI-INF-3DHP [34] are provided in the supplementary Sec. C). Among these datasets, Human3.6M is an indoor dataset while 3DPW contains complex outdoor scenes. Following [6, 50], 3DPW training sets are involved in training. The results are shown in Table 2. Our DiffMesh outperforms all previous methods across all evaluation metrics on the Human3.6M and 3DPW datasets. This superior performance validates the efficacy of our DiffMesh framework for recovering high-quality human mesh structures from video input. Specifically, without involving an additional refinement procedure, our DiffMesh surpasses the second-best method MotionBERT [64] (which already exhibits significant improvements over its predecessors) by more than 1 mm of MPJPE, PA-MPJPE, and MPVE. When incorporating HybrIK [24] as an additional refinement procedure, our DiffMesh consistently surpasses MotionBERT [64] with refinement, which significantly enhances the performance of video-based HMR.
@@
Method Year Input frames MPVE $\downarrow$ ACC-ERR $\downarrow$
Image-based SPIN [21] 2019 1 116.4 29.8
HybrIK [24] 2021 1 89.1 22.8
POTTER [61] 2023 1 87.4 23.1
Video-based TCMR [6] 2021 16 102.9 7.1
MPS-Net [50] 2022 16 99.7 7.4
GLoT [42] 2023 16 96.3 6.6
DND [25] 2022 32 88.6 7.0
DiffMesh (Ours) 16 86.4 6.1
Table 3: Performance comparison with image-based methods and video-based methods on 3DPW dataset.
To further validate the enhancement in temporal consistency achieved by DiffMesh, we conducted an evaluation of acceleration error as shown in Table 3. While some image-based methods can achieve impressive performance in terms of accuracy (low MPVE and MPJPE), they often suffer from motion jitter and temporal inconsistency issues. For instance, we observed that methods like HybrIK [24] and POTTER [61] exhibit lower MPVE and MPJPE, but concurrently, their ACC-ERR are notably higher compared to video-based HMR methods such as TCMR [6], MPS-Net [50], and GLoT [42]. Instead, our DiffMesh, which integrates motion patterns into the diffusion model and enhances temporal consistency through successive steps across frames, manifests in its superior performance. We further reduce the ACC-ERR from 6.6 to 6.1, which is about 7.6% reduction over the previous best ACC-ERR from GLoT [42]. These results underscore the practical viability of DiffMesh for real-world applications, particularly in scenarios requiring the generation of accurate and smooth human mesh sequences.
@@
HybrIK [23] 1 1 89.7 76.2 29.8 38.9 ms 38.9 ms/frame
MPS-Net [50] 16 1 99.7 84.3 7.4 31.1 ms 31.1 ms/frame
@@
Table 4: Reconstruction performance and inference time comparison on 3DPW dataset between our DiffMesh and previous video-based HMR methods with the same hardware platform (a single NVIDIA A5000 GPU with a batch size of 1 for fair comparison).
4.4 Ablation Study
@@
Comparison with the Baselines: In Table 4, we present a performance comparison among Baselines (in Section 3.2) and our DiffMesh. Baseline 1 applies the diffusion model frame by frame, achieving better MPVE and PA-MPJPE results than Baseline 2. However, it falls short in terms of ACC-ERR. When adopting the concatenation and separation strategy in Baseline 2, the ACC-ERR is improved, but with a slight downgrade in MPVE and PA-MPJPE. Neither Baseline 1 nor Baseline 2 surpass the previous method DND [25]. In contrast, our DiffMesh outperforms both Baseline 1 and Baseline 2 across all evaluation metrics. Furthermore, DiffMesh demonstrates a 12.8% reduction in ACC-ERR compared to the SOTA DND [25] even with fewer input frames (16 vs. 32). We further visualize the comparison of the ACC-ERR in Fig. 6 (“downtown_bar_00_0” of 3DPW test-set). The ACC-ERR for DND [25] is unavailable in Fig. 6 as their code has not been fully released. This highlights the superiority of DiffMesh, which leverages the consideration of human motion patterns within the diffusion model, resulting in enhanced overall performance, especially in terms of motion smoothness.
@@
Figure 6: Comparison of the acceleration errors for MPS-Net [50], two baselines, and our DiffMesh.
Inference Time Analysis: For video-based HMR, inference speed is indeed a crucial evaluation metric. While methods like TCMR [6] and MPS-Net [50] employ video input to enhance motion smoothness and temporal consistency, they often face limitations in terms of processing speed. For instance, when provided with a sequence of frames (e.g., 16 frames) as input, MPS-Net [50] can only generate the human mesh for one frame in Table 4. The average processing time of these methods exceeds 30 ms/frame, indicating a need for further improvements in computational efficiency. Our Baseline 1 also requires a long processing time to output high-quality human mesh. Our Baseline 2 reduced the processing time but the acceleration error is still worse than DND [25].
@@
Compared to these methods, DiffMesh not only achieves the best mesh recovery performance and motion smoothness, while maintaining a competitive processing time. The average time can achieve 13.9 ms/frame, which is much faster than previous SOTA methods. This combination of accuracy and efficiency makes DiffMesh a promising solution for video-based human mesh recovery applications.
Number of steps in diffusion: Following previous video-based HMR methods [20, 6, 50, 42], our DiffMesh operates with an input sequence of $f=16$ frames. As introduced in Section 3.3, we assume that human motion will reach an initial distribution (we use the mesh template [30] as this static state) after additional steps as denoted in Fig. 4. The role of initial distribution is similar to the Gaussian noise in conventional diffusion models. To verify the impact of these additional steps, we conducted experiments as shown in Table 5. When introducing additional steps, the reverse motion process starts from the initial distribution (this static state is represented by the template mesh [30]). The results demonstrate that introducing additional steps can enhance overall performance. The optimal balance between efficiency and performance is achieved when the number of additional steps is set to 15 (the total number of steps $N$ is 30).
input frames
steps for output
sequence
16 15 0 15 88.5 77.4 6.5
16 15 5 20 88.0 77.1 6.5
16 15 15 30 86.4 75.7 6.1
16 15 25 40 87.1 75.6 6.2
Table 5: Performance with the different number of steps on 3DPW.
Due to space limitation, more ablation studies are available in the supplementary Sec. C, covering topics such as the number of input frames, the two-stream transformer network (Fig. 5), and the auxiliary loss.
@@
Figure 7: The in-the-wild visual comparison between recent GLoT [42] with our DiffMesh. The circles highlight locations where DiffMesh is more accurate than GLoT. More examples are provided in the supplementary Sec. D and in demo videos.
@@ 4.5 In-the-Wild Visualization
To demonstrate the practical performance of our DiffMesh, we present the qualitative comparison between DiffMesh and a recent GLoT [42] on in-the-wild video in Fig. 7. Although achieving excellent performance in terms of mesh recovery performance on common datasets, GLoT [42] suffers from temporal inconsistency when applied for in-the-wild video. As shown in Fig. 7, GLoT [42] fails to generate the accurate mesh. In contrast, our DiffMesh enhances motion smoothness and reconstruction accuracy through our proposed motion-aware diffusion-like framework.
@@ 5 Conclusion
In this paper, we present a novel diffusion-like framework (DiffMesh) for human mesh recovery from a video. DiffMesh innovatively connects diffusion models with human motion, resulting in the efficient generation of highly precise and seamlessly smooth output mesh sequences. Extensive experiments are conducted on Human3.6M and 3DPW datasets showing impressive performance compared to the SOTA video-based methods.
@@ B.1 Related Work
The majority of methods  [16, 21, 5, 60, 62, 53, 61, 26] for HMR rely on a parametric human model, such as SMPL [31], to reconstruct the mesh by estimating pose and shape parameters. As a fundamental HMR work, SPIN  [21] combines regression and optimization in a loop, where the regressed output serves as better initialization for optimization (SMPLify). METRO [30] is the first transformer-based method that models vertex-vertex and vertex-joint interaction using a transformer encoder after extracting image features with a CNN backbone. HybrIK [24] and HybrIK-X [27] present novel hybrid inverse kinematics approaches that transform 3D joints to body-part rotations via twist-and-swing decomposition. Lin et al. [29] propose a one-stage pipeline for 3D whole-body (body, hands, and face) mesh recovery. Foo et al. [11] first introduce a diffusion-based approach for recovering human mesh from a single image. The recovered human mesh is obtained by the reverse diffusion process. However, when applied to video sequences, these image-based methods suffer from severe motion jitter due to frame-by-frame reconstruction, making them unsuitable for practical use.
@@ B.2 Datasets
PoseTrack [1] is a 2D benchmark designed for multi-person pose estimation and tracking in videos. This dataset comprises 1.3K videos and 46K annotated frames in total, captured at varying fps around 25 fps. There are 792 videos used for the official train set, which includes 2D pose annotations for 30 frames located in the middle of each video.
@@ B.3 Loss Function
This auxiliary loss contributes to the refinement of our transformer-based diffusion model during the training process. Thus, the overall loss for our DiffMesh is the sum of the $\mathcal{L}_{HMR}$ and $\mathcal{L}_{aux}$:
@@
where $w_{4}=0.01$.
@@
B.4 More Details about the Architecture
Conditional features generation block: Our chosen backbone to extract features for both our proposed method and the baselines is DSTformer [64]. After extracting features from each frame $b_{i}$, where $i\in\left\{1,\dots f\right\}$, using the backbone, our goal is to generate $N$ conditional features to be utilized during the reverse process. To achieve this, we pad additional $N-f$ zero features, $b_{f+1},\cdots,b_{N}$. Then, we combine them with the existing features, creating new features $b\in\mathbb{R}^{N\times D}$, where $D$ represents the embedded dimension. Subsequently, we apply a transformer block [9] to model these features and return the required conditional features denoted as $c\in\mathbb{R}^{N\times D}$.

### v4

@@ 4 Experiments
Video-based Method Venue Backbone MPVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ ACC-ERR $\downarrow$ MPVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ ACC-ERR $\downarrow$
Fixed backbone
VIBE [22] CVPR 2020 ResNet50[14] - 65.6 41.4 - - 91.9 57.6 25.4
TCMR [5] CVPR 2021 ResNet50[14] - 62.3 41.1 5.3 102.9 86.5 52.7 6.8
MPS-Net [54] CVPR 2022 ResNet50[14] - 69.4 47.4 3.6 99.7 84.3 52.1 7.4
GLoT [45] CVPR 2023 ResNet50[14] - 67.0 46.3 3.6 96.3 80.7 50.6 6.6
STAF [57] TCSVT 2024 ResNet50[14] - - - - 95.3 80.6 48.0 8.2
DiffMesh (ours) ResNet50[14] - 65.3 41.9 3.3 94.0 77.2 48.5 6.3
Other backbone
MAED [52] ICCV 2021 ResNet50[14]+ViT[9] 84.1 60.4 38.3 - 93.3 79.0 45.7 17.6
MotionBERT [70] ICCV 2023 DSTformer [70] 65.5 53.8 34.9 - 88.1 76.9 47.2 -
DiffMesh (Ours) DSTformer [70] 64.2 52.5 33.5 5.5 86.4 75.7 45.6 6.1
@@
DND [26] $\dagger$ ECCV 2022 HybrIK [28] - 52.5 35.5 6.1 88.6 73.7 42.7 7.0
MotionBERT [70] $\dagger$ ICCV 2023 DSTformer [70] 52.6 43.1 27.8 - 79.4 68.8 40.6 -
DiffMesh (Ours) $\dagger$ DSTformer [70] 52.1 42.4 27.7 5.7 78.4 67.8 40.1 6.3
Table 2: Performance comparison with SOTA video-based methods on Human3.6M and 3DPW datasets. The symbol $\dagger$” denotes the HybrIK [28] is applied for the refinement. 3DPW Training set is used during training.
@@ 4.1 Implementation Details
For fair comparisons, we adhere to the standard implementation details commonly employed in the video-based HMR task, consistent with prior works such as [22, 5, 54, 70]. We set the length of the input sequence $f$ to 16. The number of steps $N=30$. We adopt the same variance schedules as DiffPose [13] during our training and testing. Our approach is implemented using PyTorch [39] on a single NVIDIA A5000 GPU. The optimization of weights is conducted using the Adam optimizer [21], with an initial learning rate of $5e^{-6}$ and weight decay set to 0.98. During the training phase, we utilize a batch size of 64, and the training process spans 60 epochs. More detailed information about the conditional feature generation part is included in the supplementary Sec. 2.
@@ 4.3 Comparison with State-of-the-art Methods
We compare our DiffMesh with previous SOTA video-based methods on Human3.6M and 3DPW datasets (results on MPI-INF-3DHP [36] are provided in the supplementary Sec. 3). Among these datasets, Human3.6M is an indoor dataset while 3DPW contains complex outdoor scenes. Following [5, 54], 3DPW training sets are involved in training. The results are shown in Table 2. Our DiffMesh outperforms all previous methods across all evaluation metrics on the Human3.6M and 3DPW datasets. This superior performance validates the efficacy of our DiffMesh framework for recovering high-quality human mesh structures from video input. Specifically, Without involving an additional refinement procedure, our DiffMesh surpasses the second-best method MotionBERT [70] (which already exhibits significant improvements over its predecessors) by more than 1 mm of MPJPE, PA-MPJPE, and MPVE. When incorporating HybrIK [28] as an additional refinement procedure, our DiffMesh consistently surpasses MotionBERT [70] with refinement, which significantly enhances the performance of video-based HMR.
@@
Video-based Method Backbone MPVE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ ACC-ERR $\downarrow$
VIBE [22] Fixed ResNet50[14] 113.4 93.5 56.5 27.1
TCMR [5] Fixed ResNet50[14] 111.5 95.0 55.8 7.0
MPS-Net [54] Fixed ResNet50[14] 109.6 91.6 54.0 7.5
GLoT [45] Fixed ResNet50[14] 107.8 89.9 53.5 6.7
DiffMesh (ours) Fixed ResNet50[14] 105.9 88.7 53.0 6.5
Table 3: Performance comparison with SOTA video-based methods on 3DPW without using 3DPW training set during training.
Following previous methods [5, 54, 45], we report the results on the 3DPW [51] without using the 3DPW training set in Table. 3. To ensure a fair comparison, we also fixed ResNet50 [14] as our backbone. Our DiffMesh consistently surpasses those methods across all evaluation metrics.
4.4 Ablation Study
@@
Image-based
I2L-MeshNet [38] 1 1 110.1 93.2 30.9 41.1 ms 41.1 ms/frame
HybrIK [29] 1 1 89.7 76.2 22.8 38.9 ms 38.9 ms/frame
POTTER [66] 1 1 87.4 75.0 23.1 42.7 ms 42.7 ms/frame
Video-based
@@
Table 4: Reconstruction performance and inference time comparison on 3DPW dataset between our DiffMesh and previous video-based HMR methods with the same hardware platform. A single NVIDIA A5000 GPU is used with a batch size of 1 (the input of [1, num of frames, 224, 224, 3]) for fair comparison.
@@
Comparison with the Baselines: In Table 4, we present a performance comparison among Baselines (in Section 3.2) and our DiffMesh. Baseline 1 applies the diffusion model frame by frame, achieving better MPVE and PA-MPJPE results than Baseline 2. However, it falls short in terms of ACC-ERR. When adopting the concatenation and separation strategy in Baseline 2, the ACC-ERR is improved, but with a slight downgrade in MPVE and PA-MPJPE. Neither Baseline 1 nor Baseline 2 surpass the previous method DND [26]. In contrast, our DiffMesh outperforms both Baseline 1 and Baseline 2 across all evaluation metrics.
Temporal consistency and motion smoothness: To further validate the enhancement in temporal consistency achieved by DiffMesh, we conducted an evaluation of acceleration error as shown in Table 4. While some image-based methods can achieve impressive performance in terms of accuracy (low MPVE and MPJPE), they often suffer from motion jitter and temporal inconsistency issues. For instance, we observed that methods like HybrIK [28] and POTTER [66] exhibit lower MPVE and MPJPE, but concurrently, their ACC-ERR are notably higher compared to video-based HMR methods . Instead, our DiffMesh, which integrates motion patterns into the diffusion model and enhances temporal consistency through successive steps across frames, manifests in its superior performance. DiffMesh demonstrates a 12.8% reduction in ACC-ERR compared to the SOTA DND [26] even with fewer input frames (16 vs. 32). We further visualize the comparison of the ACC-ERR in Fig. 7 (“courtyard_basketball_01’ of 3DPW dataset). The ACC-ERR for DND [26] is unavailable in Fig. 7 as their code has not been fully released. This highlights the superiority of DiffMesh, which leverages the consideration of human motion patterns within the diffusion model, resulting in enhanced overall performance, especially in terms of motion smoothness.
@@
Figure 7: Acceleration error Comparison of the ‘courtyard_basketball_01’ sequence for TCMR [5], two baselines, and our DiffMesh.
@@
Inference time analysis: A prevalent consideration regarding diffusion-based methods pertains to their inference speed, which is also a crucial evaluation metric besides accuracy. Our Baseline 1 requires a long processing time to output high-quality human mesh. Our Baseline 2 reduced the processing time but the acceleration error is still worse than DND [26]. Compared to these methods, DiffMesh not only achieves the best mesh recovery performance and motion smoothness, while maintaining a competitive processing time. The average time can achieve 13.9 ms/frame, which is much faster than previous SOTA methods (More detailed comparisons with previous methods are provided in the supplementary Sec. 3 when considering some methods can extract all features by their backbone and then utilize batch processing to accelerate the inference speed). This combination of accuracy and efficiency makes DiffMesh a promising diffusion solution for video-based human mesh recovery applications.
@@
Figure 8: The in-the-wild visual comparison between recent GLoT [45] with our DiffMesh. The circles highlight locations where DiffMesh is more accurate than GLoT. More examples are provided in the supplementary Sec. 4 and in demo videos.
@@ 4.5 In-the-Wild Visualization
To demonstrate the practical performance of DiffMesh, we present the qualitative comparison between DiffMesh and a recent GLoT [45] on in-the-wild video in Fig. 8. Although achieving good performance in terms of mesh recovery performance on common datasets, GLoT [45] suffers from temporal inconsistency when applied for in-the-wild video. As shown in Fig. 8, GLoT [45] fails to generate the accurate mesh. In contrast, our DiffMesh enhances motion smoothness and reconstruction accuracy through our proposed motion-aware diffusion framework.
@@ 5 Conclusion
In this paper, we present a novel diffusion framework (DiffMesh) for human mesh recovery from a video. DiffMesh innovatively connects diffusion models with human motion, resulting in the efficient generation of highly precise and seamlessly smooth output mesh sequences. Extensive experiments are conducted on Human3.6M and 3DPW datasets showing impressive performance compared to the SOTA video-based methods.
@@ B.1 Related Work
The majority of methods  [17, 23, 6, 67, 68, 58, 66, 24] for HMR rely on a parametric human model, such as SMPL [33], to reconstruct the mesh by estimating pose and shape parameters. As a fundamental HMR work, SPIN  [23] combines regression and optimization in a loop, where the regressed output serves as better initialization for optimization (SMPLify). METRO [32] is the first transformer-based method that models vertex-vertex and vertex-joint interaction using a transformer encoder after extracting image features with a CNN backbone. HybrIK [28] and HybrIK-X [25] present novel hybrid inverse kinematics approaches that transform 3D joints to body-part rotations via twist-and-swing decomposition. Lin et al. [31] propose a one-stage pipeline for 3D whole-body (body, hands, and face) mesh recovery. PyMAF [63] and its extension work PyMAF-x [62] capitalize on a feature pyramid to rectify predicted parameters by aligning meshes with images, extracting mesh-aligned evidence from finer-resolution features. CLIFF [30] enhances holistic feature representation by incorporating bounding box information into cropped-image features. It employs a 2D reprojection loss considering the full frame and leverages global-location aware supervision to directly predict global rotation and more accurately articulated poses. ReFit [53] proposes a feedback-update loop reminiscent of solving inverse problems via optimization, iteratively reprojections keypoints from the human model to feature maps for feedback, and utilizes a recurrent-based updater to refine the model’s fit to the image. HMR2.0 [12] develops a system that can simultaneously reconstruct and track humans from video, but only reports the frame-based results for the HMR task without considering temporal information. Foo et al. [11] first introduce a diffusion-based approach for recovering human mesh from a single image. The recovered human mesh is obtained by the reverse diffusion process. However, when applied to video sequences, these image-based methods suffer from severe motion jitter due to frame-by-frame reconstruction, making them unsuitable for practical use.
@@ B.2 Datasets
PoseTrack [1] is a 2D benchmark designed for multi-person pose estimation and tracking in videos. This dataset comprises 1.3K videos and 46K annotated frames, captured at varying fps around 25 fps. There are 792 videos used for the official train set, which includes 2D pose annotations for 30 frames located in the middle of each video.
@@ B.3 Loss Function
This auxiliary loss contributes to the refinement of our transformer-based diffusion model during the training process. Thus, the overall loss for our DiffMesh is the sum of the $\mathcal{L}_{HMR}$, $\mathcal{L}_{aux}$ and $\mathcal{L}_{diffusion}$:
@@
where $w_{4}=0.01$ and $w_{5}=0.01$.
B.4 More Details about the Architecture
@@
Conditional features generation block: Our chosen backbone to extract features for both our proposed method and the baselines is ResNet-50 [14] or DSTformer [70]. After extracting features from each frame $b_{i}$, where $i\in\left\{1,\dots f\right\}$, using the backbone, our goal is to generate $N$ conditional features to be utilized during the reverse process. To achieve this, we pad additional $N-f$ zero features, $b_{f+1},\cdots,b_{N}$. Then, we combine them with the existing features, creating new features $b\in\mathbb{R}^{N\times D}$, where $D$ represents the embedded dimension. Subsequently, we apply a transformer block [9] to model these features and return the required conditional features denoted as $c\in\mathbb{R}^{N\times D}$.